A ball rolls off the edge of a horizontal platform 2meters high and lands 2.5 meters from the edge of the platform. Find (a) the time of its flight and; (b) its initial horizontal velocity


1
Expert's answer
2022-05-08T14:29:41-0400

a)

"t=\sqrt\frac{2h}g=0.64~s."

b)

"v=\frac st=3.9~\frac ms."